OpenStudy (anonymous):

A ceiling fan has 19 -inch blades (so the radius of the circular fan is 19 inches). Suppose the fan turns at a rate of 41 revolutions per minute. (a) Find the angular speed of the fan in radians per minute. (b) Find the linear speed of the tip of a blade in feet per second. Do not round any intermediate computations, and round your answer to the nearest whole number.

OpenStudy (aum):

1 revolution is \(2\pi\) radians. 41 revolutions per minute is \(41 * 2\pi\) radians / min = \(82\pi\) radians / min.

OpenStudy (aum):

b) \(v = r\omega\) r = 19 in. = 19/12 ft. \(\omega = 82\pi\) radians per min = \(82\pi/60\) radians per second. Find v.